# Probabilistically Plausible Counterfactual Explanations with Normalizing Flows

### v1

@@ Abstract
We present PPCEF, a novel method for generating probabilistically plausible counterfactual explanations (CFs). PPCEF advances beyond existing methods by combining a probabilistic formulation that leverages the data distribution with the optimization of plausibility within a unified framework. Compared to reference approaches, our method enforces plausibility by directly optimizing the explicit density function without assuming a particular family of parametrized distributions. This ensures CFs are not only valid (i.e., achieve class change) but also align with the underlying data’s probability density. For that purpose, our approach leverages normalizing flows as powerful density estimators to capture the complex high-dimensional data distribution. Furthermore, we introduce a novel loss that balances the trade-off between achieving class change and maintaining closeness to the original instance while also incorporating a probabilistic plausibility term. PPCEF’s unconstrained formulation allows for efficient gradient-based optimization with batch processing, leading to orders of magnitude faster computation compared to prior methods. Moreover, the unconstrained formulation of PPCEF allows for the seamless integration of future constraints tailored to specific counterfactual properties. Finally, extensive evaluations demonstrate PPCEF’s superiority in generating high-quality, probabilistically plausible counterfactual explanations in high-dimensional tabular settings. This makes PPCEF a powerful tool for not only interpreting complex machine learning models but also for improving fairness, accountability, and trust in AI systems.
@@ 1 Introduction
Counterfactual explanations (briefly counterfactuals, and abbreviated as CF) are one particular type of such explanations of black box model predictions that provide information about how feature values of an example should be changed to obtain a more desired prediction of the model (i.e., to change its target decision). On the one hand, by interacting with the model using counterfactuals, the user can better understand how the system works by exploring "what would have happened if…" scenarios. On the other hand, a good counterfactual provides a practical recommendation to the user about what changes are needed in order to achieve the desired outcome.
@@
Up to now, several algorithms for generating counterfactual explanations have been introduced. They are based on different principles, and for comprehensive surveys, see, e.g., [10, 28]. Depending on the specific method, some properties of counterfactuals are expected to be met, such as validity of the decision change, proximity to the input instance, sparsity of recommended changes, their actionability, and plausibility of locating the counterfactual within a high-density region of the data, ensuring that the proposed counterfactuals are realistic and feasible within the context of the observed data distribution.
@@
In previous works, plausibility has often been verified by simple -neighbourhood analysis of the counterfactual with respect to the original data [25, 14, 26]. Some other approaches [3] model the conditional density in the target class and try to find the counterfactual example with the density value above the given threshold. Although the problem is quite well mathematically defined, the current methods apply simple approaches like kernel density estimators or a mixture of Gaussians to model conditional distributions that are difficult to apply for high-dimensional data. Moreover, the problem of estimating valid and plausible counterfactuals is defined as a complex constrained optimization problem with strict convexity assumptions [3]. Finally, the currently proposed methods, while providing valid counterfactuals, struggle to consistently produce observations that fulfill the plausibility criteria.
@@ 3 Background
This approach’s crucial aspect is finding the proper model to represent the conditional density function . Typically, kernel density estimators (KDEs) are used to model conditional densities, but the use of non-linear kernels results in highly non-convex optimization problem formulation. Gaussian Mixture Model (GMM) can be applied alternatively, but convexity constraints are still not satisfied. To facilitate the desired optimization process, the authors of Artelt and Hammer [3] propose to approximate the density value using a component-wise maximum of GMM components:
@@ 4.1 Unconstrained Probabilistically Plausible Counterfactual Explanations
We consider a binary classification problem, . However, our considerations can be easily extended to the multiclass case. Further, we consider a discriminative differentiable model (i.e., Logistic Regression or MLP) and reformulate the validity constrain as , where , practically represented as small enough value close to .
@@
The loss component controls the validity constrain and is defined as follows:
@@ 4.2 Probabistically Plausible Counterfactual Explanations via Normalizing Flow-based Density Estimation
KDE or GMMs can be used to model the conditional distributions. However, those models have limited modeling capabilities due to the parametrized (usually Gaussian) form of or the inability to model high-dimensional data (KDE). Therefore, in this work, we postulate the use of a conditional normalizing flow model [23] to estimate the density for the joined distribution of the attributes for each class.
@@ 5 Experiments
In this section, we aim to demonstrate and validate our counterfactual explanation method through a series of experiments. Initially, we illustrate our method’s intuition with the Moons dataset and Logistic Regression model. Next, we compare our approach against the only reference method in a probabilistically plausible CFs area - Artelt and Hammer [3], as well as other established CF methods. This comparison focuses on the impact of plausibility on proximity metrics and time efficiency. Lastly, we conduct broader comparisons using other classifier models: Logistic Regression (LR), Multilayer Perceptron (MLP), and Neural Oblivious Decision Ensembles (NODE). The code for these experiments is included in the Supplementary Materials and will be publicly released on GitHub after the review process.
@@
To evaluate PPCEF’s effectiveness, we conducted experiments on seven numerical-only tabular datasets. Four datasets (Law, Heloc, Moons, and Audit) represent binary classification problems, whereas the first two datasets (Law and Heloc) are commonly used benchmarks for counterfactual explanation tasks. The remaining three datasets (Blobs, Digits, and Wine) address multiclass classification problems. Detailed descriptions of these datasets are available in the Appendix B. Overall, they represent broad diversity in sample sizes, number of variables, and number of classes. For preprocessing purposes, we implemented two key steps to prepare the datasets. First, we addressed class imbalance by downsampling the majority class to match the size of the minority class. Second, we normalized all features across the datasets to a range, enabling consistent scale and comparability among features. Thirdly, to ensure robust method evaluation, we employed stratified 5-fold cross-validation on each dataset. Finally, for clarity, the main manuscript reports average values, while the appendix includes standard deviation for detailed analysis.
@@
For the experiments, we include Logistic Regression (LR), 3-layer Multilayer Perceptron (MLP), and Neural Oblivious Decision Ensemble (NODE) [21] catering to both linear and non-linear scenarios. LR aligns with linear assumptions prevalent in some baseline methods, MLP allows for the assessment of behaviors in non-linear model contexts, and NODE stands as an example of a complex ensemble of neural decision trees. This triple-model approach facilitates a thorough evaluation across varied model complexities. Crucially, all models are differentiable, which is essential in the context of our method.
@@
For every combination of the classification model and dataset, we trained both the classification model and a Normalizing Flow as the density estimator, following the approach detailed in Section 4.2. We opted for the Masked Autoregressive Flow (MAF) architecture [19] as our choice for the Normalizing Flow. This decision was based on experimental findings indicating MAF’s superior performance in accurately fitting data distributions. For a deeper analysis of these results, including in-depth model performance metrics like accuracy, please refer to the Appendix. See Section C for a detailed exploration and Tab. 9 for specific performance figures. The final step involved generating counterfactual explanations for the entire set of test samples.
@@ 5.2 Probabilistically Plausibile Counterfactual Explanations Methods Comparison
The results are presented in Tab. 1. Firstly, we can observe that our method always returns the results that are probabilistically plausible. That is not the case for Artelt’s method, which struggles in high-dimensional datasets like Heloc (23 dimensions) or Digits (64 dimensions), doesn’t support non-linear classifiers like MLPs, and wasn’t able to consistently fulfill the probabilistic plausibility criterion. Secondly, in terms of distances, Artelt’s method returns better results, which is expected due to the trade-off between distance and plausibility, i.e., the more plausible observations, the farther away they usually are. However, the results are not significantly worse, especially in terms of L2 distance, meaning PPCEF can balance both desired properties of counterfactuals. Thirdly, the log density values of the observations produced by PPCEF method are significantly better. Fourthly, our analysis using Local Outlier Factor (LOF) and Isolation Forest (IsoForest) metrics indicates that our methods generate inliers (except for Audit and Heloc, where almost all methods struggle to obtain reasonable values of LOF), whereas Artelt’s method underperforms and can sometimes result in outliers. Fifthly, our method turned out to be significantly faster, with the speed up around x2-10 on relatively small datasets. Finally, our method was almost always able to produce valid counterfactual explanations for MLP and NODE, contrary to Artelt (see results in Tab. 2 and detailed results in Tab. 6 and 7 in Appendix). It’s worth mentioning that PPCEF almost always returned probabilistically plausible observations, which, in case of non-valid observations, might still be valuable insight for the final user, contrary to the lack of a response at all.
@@ 5.3 Counterfactual Explanations Methods Comparison
Results presented in Tab. 1 and Tab. 2 indicate that our model excels in validity, plausibility (considering both probabilistic formulation and outlier metrics), and processing times while maintaining reasonable distances compared to competing approaches across all datasets and classification methods. Specifically, Tab. 2 presents the evaluation results for two selected high-dimensional datasets (one for binary classification problem and one for multiclass problem) using two advanced classifiers, demonstrating that our method consistently produces valid results not only with a shallow model, such as LR but also with deeper models, including MLP and NODE. In contrast, the majority of existing methods encounter difficulties in producing valid counterfactual explanations for the Multilayer Perceptron. We conducted a comprehensive evaluation utilizing all methods and datasets mentioned earlier, applying three different classifiers. Detailed outcomes are presented in Appendix A. Particularly, results for Logistic Regression are shown in Tab. 5, while findings for the MLP and NODE classifiers are detailed in Tab. 6 and 7, respectively.
@@ 6.2 Regularization Hyperparameter Analysis
To evaluate the impact of the regularization hyperparameter on the fulfillment of validity and probabilistic plausibility conditions, we conducted a focused hyperparameter sensitivity analysis. While theoretically should extend to infinity, practical considerations necessitate setting a feasible value. Our objective is to identify an optimal that not only guarantees condition fulfillment but also to understand its influence on other metrics. Experiments were carried out on the Moons and Law datasets, exploring values within the set . The results in Tab. 4 indicate that moderate values of , like or , deliver satisfactory outcomes, while values around or more almost invariably guarantee the fulfillment of the conditions, leading us to adopt the value of for all preceding experiments. This experiment also confirms the expected trade-off: higher strictness in counterfactual conditions results in decreased proximity metrics, necessitating larger deviations from the original data point.
@@ Appendix B Datasets
In Tab. 9, we provide detailed descriptions of the datasets utilized in our study: Moons, Law111https://www.kaggle.com/datasets/danofer/law-school-admissions-bar-passage, Audit222https://archive.ics.uci.edu/dataset/475/audit+data, Heloc333https://community.fico.com/s/explainable-machine-learning-challenge, Wine444https://archive.ics.uci.edu/dataset/109/wine, Blobs and Digits555https://archive.ics.uci.edu/dataset/80/optical+recognition+of+handwritten+digits. The Moons dataset is an artificially generated set comprising two interleaving half-circles. It includes a standard deviation of Gaussian noise set at 0.01. The Law dataset originates from the Law School Admissions Council (LSAC) and is referred to in the literature as the Law School Admissions dataset [31]. For our analysis, we selected three features that exhibit the highest correlation with the target variable: entrance exam scores (LSAT), grade-point average (GPA), and first-year average grade (FYA). The Audit dataset, which encompasses comprehensive one-year non-confidential data of firms in the years 2015 to 2016, is collected from the Auditor Office of India to build a predictor for classifying suspicious firms. The Heloc dataset, initially utilized in the ’FICO xML Challenge’, consists of Home Equity Line of Credit (HELOC) applications submitted by real homeowners. This dataset comprises various numeric features that encapsulate information from the applicant’s credit report. The primary objective is to predict whether the applicant will repay their HELOC account within a two-year period. This prediction is instrumental in determining the applicant’s qualification for a line of credit. The Wine dataset comprises chemical analysis results for wines originating from the same region in Italy, produced from three distinct cultivars. This analysis quantified 13 different constituents present in each of the three wine varieties. The Blobs dataset is an artificially generated isotropic Gaussian blobs, characterized by equal variance. The Digits dataset is utilized for the optical recognition of handwritten digits. It consists of 32x32 bitmap images that are segmented into non-overlapping 4x4 blocks. Within each block, the count of ’on’ pixels is recorded, resulting in an 8x8 input matrix. Each element of this matrix is an integer between 0 and 16.
@@ Appendix D Implementation details
In the implementation of our experiments, we utilized Python as the primary programming language. The core of our computational framework was PyTorch, a popular open-source machine learning library. A key feature of our implementation was the gradient optimization approach, designed to be executed in batches. This approach was particularly effective, allowing us to process entire test sets in a single batch. Our experiments were conducted on an M1 Apple Silicon CPU paired with 16GB of RAM. This hardware setup was more than sufficient for our experiment needs, providing enough capacity for the computational demands of our algorithm while ensuring fast processing speeds.

### v2

@@ Abstract
We present PPCEF, a novel method for generating probabilistically plausible counterfactual explanations (CFs). PPCEF advances beyond existing methods by combining a probabilistic formulation that leverages the data distribution with the optimization of plausibility within a unified framework. Compared to reference approaches, our method enforces plausibility by directly optimizing the explicit density function without assuming a particular family of parametrized distributions. This ensures CFs are not only valid (i.e., achieve class change) but also align with the underlying data’s probability density. For that purpose, our approach leverages normalizing flows as powerful density estimators to capture the complex high-dimensional data distribution. Furthermore, we introduce a novel loss function that balances the trade-off between achieving class change and maintaining closeness to the original instance while also incorporating a probabilistic plausibility term. PPCEF’s unconstrained formulation allows for an efficient gradient-based optimization with batch processing, leading to orders of magnitude faster computation compared to prior methods. Moreover, the unconstrained formulation of PPCEF allows for the seamless integration of future constraints tailored to specific counterfactual properties. Finally, extensive evaluations demonstrate PPCEF’s superiority in generating high-quality, probabilistically plausible counterfactual explanations in high-dimensional tabular settings.
@@ 1 Introduction
Counterfactual explanations (briefly counterfactuals, and abbreviated as CF) are one particular type of such explanations of black box model predictions that provide information about how feature values of an example should be changed to obtain a more desired prediction of the model (i.e., to change its target decision) [30]. On the one hand, by interacting with the model using counterfactuals, the user can better understand how the system works by exploring "what would have happened if…" scenarios. On the other hand, a good counterfactual provides a practical recommendation to the user about what changes are needed in order to achieve the desired outcome.
@@
Up to now, several algorithms for generating counterfactual explanations have been introduced. They are based on different principles, and for comprehensive surveys, see, e.g., [10, 30]. Depending on the specific method, some properties of counterfactuals are expected to be met, such as validity of the decision change, proximity to the input instance, sparsity of recommended changes, their actionability, i.e., the counterfactual should not modify immutable features or violate monotonic constraints, and plausibility of locating the counterfactual within a high-density region of the data, ensuring that the proposed counterfactuals are realistic and feasible within the context of the observed data distribution.
@@
In previous works, plausibility has often been verified by simple -neighbourhood analysis of the counterfactual with respect to the original data [26, 14, 27]. Few other approaches [3] model the conditional density in the target class and try to find the counterfactual example with the density value above the given threshold. Although the problem is quite well mathematically defined, the current methods apply simple approaches like kernel density estimators or a mixture of Gaussians to model conditional distributions that are difficult to apply for high-dimensional data. Moreover, the problem of estimating valid and plausible counterfactuals is defined as a complex constrained optimization problem with strict convexity assumptions [3]. Finally, the currently proposed methods, while providing valid counterfactuals, struggle to consistently produce observations that fulfill the plausibility criteria.
@@ 3 Background
This approach’s crucial aspect is finding the proper model to represent the conditional density function . Typically, kernel density estimators (KDEs) are used to model conditional densities, but the use of non-linear kernels results in the highly non-convex optimization problem formulation. Gaussian Mixture Model (GMM) can be applied alternatively, but convexity constraints are still not satisfied. To facilitate the desired optimization process, the authors of Artelt and Hammer [3] propose to approximate the density value using a component-wise maximum of GMM components:
@@ 4.1 Unconstrained Probabilistically Plausible Counterfactual Explanations
We consider a binary classification problem, . However, our considerations can be easily extended to the multiclass case. Further, we consider a discriminative differentiable model (e.g., Logistic Regression or MLP) and reformulate the validity constraint as , where , practically represented as small enough value close to .
@@
The loss component controls the validity constraint and is defined as follows:
@@ 4.2 Probabistically Plausible Counterfactual Explanations via Normalizing Flow-based Density Estimation
KDE or GMMs can be used to model the conditional distributions. However, those models have limited modeling capabilities due to the parametrized (usually Gaussian) form of or the inability to model high-dimensional data (KDE). Therefore, in this work, we postulate the use of a conditional normalizing flow model [24] to estimate the density for the joint distribution of the attributes for each class.
@@ 5 Experiments
In this section, we aim to demonstrate and validate our counterfactual explanation method through a series of experiments. Initially, we illustrate our method’s intuition with the Moons dataset and Logistic Regression model. Next, we compare our approach against the only reference method in a probabilistically plausible CFs area - Artelt and Hammer [3], as well as other established CF methods. This comparison focuses on the impact of plausibility on proximity metrics and time efficiency. Lastly, we conduct broader comparisons using other classifier models: Logistic Regression (LR), Multilayer Perceptron (MLP), and Neural Oblivious Decision Ensembles (NODE) [22]. The code for these experiments is publicly released on GitHub111https://github.com/ofurman/counterfactuals.
@@
To evaluate PPCEF’s effectiveness, we conducted experiments on seven numerical-only tabular datasets. Four datasets (Law, Heloc, Moons, and Audit) represent binary classification problems, whereas the first two datasets (Law and Heloc) are commonly used benchmarks for counterfactual explanation tasks. The remaining three datasets (Blobs, Digits, and Wine) address multiclass classification problems. Detailed descriptions of these datasets are available in the Appendix B [33]. Overall, they represent broad diversity in sample sizes (up to approximately 10.000), number of variables (up to 64), and number of classes (up to 10). For preprocessing purposes, we implemented two key steps to prepare the datasets. First, we addressed class imbalance by downsampling the majority class to match the size of the minority class. Second, we normalized all features across the datasets to a range, enabling consistent scale and comparability among features. Thirdly, to ensure robust method evaluation, we employed stratified 5-fold cross-validation on each dataset. Finally, for clarity, the main manuscript reports average values, while the appendix [33] includes standard deviation for detailed analysis.
@@
For the experiments, we include Logistic Regression (LR), 3-layer Multilayer Perceptron (MLP), and Neural Oblivious Decision Ensemble (NODE) catering to both linear and non-linear scenarios. LR aligns with linear assumptions prevalent in some baseline methods, MLP allows for the assessment of behaviors in non-linear model contexts, and NODE stands as an example of a complex ensemble of neural decision trees. This triple-model approach facilitates a thorough evaluation across varied model complexities. Crucially, all models are differentiable, which is essential in the context of our method.
@@
For every combination of the classification model and dataset, we trained both the classification model and a Normalizing Flow as the density estimator, following the approach detailed in Section 4.2. We opted for the Masked Autoregressive Flow (MAF) architecture [19] as our choice for the Normalizing Flow. This decision was based on experimental findings indicating MAF’s superior performance in accurately fitting data distributions. For a deeper analysis of these results, including in-depth model performance metrics like accuracy, please refer to the Appendix [33]. See Section C for a detailed exploration and Tab. 9 for specific performance figures. The final step involved generating counterfactual explanations for the entire set of test samples.
@@ 5.2 Probabilistically Plausibile Counterfactual Explanations Methods Comparison
The results are presented in Tab. 1. Firstly, we can observe that our method always returns the results that are probabilistically plausible. That is not the case for Artelt’s method, which struggles in high-dimensional datasets like Heloc (23 dimensions) or Digits (64 dimensions), doesn’t support non-linear classifiers like MLPs, and wasn’t able to consistently fulfill the probabilistic plausibility criterion. Secondly, in terms of distances, Artelt’s method returns better results, which is expected due to the trade-off between distance and plausibility, i.e., the more plausible observations, the farther away they usually are. However, the results are not clearly worse, especially in terms of L2 distance, meaning PPCEF can balance both desired properties of counterfactuals. Thirdly, the log density values of the observations produced by PPCEF method are significantly better. Fourthly, our analysis using Local Outlier Factor (LOF) and Isolation Forest (IsoForest) metrics indicates that our methods generate inliers (except for Audit and Heloc, where almost all methods struggle to obtain reasonable values of LOF), whereas Artelt’s method underperforms and can sometimes result in outliers. Fifthly, our method turned out to be significantly faster, with the speed up around x2-10 on relatively small datasets. Finally, our method was almost always able to produce valid counterfactual explanations for MLP and NODE, contrary to Artelt (see results in Tab. 2 and detailed results in Tab. 6 and 7 in Appendix [33]). It’s worth mentioning that PPCEF almost always returned probabilistically plausible observations, which, in case of non-valid observations, might still be valuable insight for the final user, contrary to the lack of a response at all.
@@ 5.3 Counterfactual Explanations Methods Comparison
Results presented in Tab. 1 and Tab. 2 indicate that our model excels in validity, plausibility (considering both probabilistic formulation and outlier metrics), and processing times while maintaining reasonable distances compared to competing approaches across all datasets and classification methods. Specifically, Tab. 2 presents the evaluation results for two selected high-dimensional datasets (one for binary classification problem and one for multiclass problem) using two advanced classifiers, demonstrating that our method consistently produces valid results not only with a shallow model, such as LR but also with deeper models, including MLP and NODE. In contrast, the majority of existing methods encounter difficulties in producing valid counterfactual explanations for the Multilayer Perceptron. We conducted a comprehensive evaluation utilizing all methods and datasets mentioned earlier, applying three different classifiers. Detailed outcomes are presented in Appendix A [33]. Particularly, results for Logistic Regression are shown in Tab. 5, while findings for the MLP and NODE classifiers are detailed in Tab. 6 and 7, respectively.
@@ 6.2 Regularization Hyperparameter Analysis
To evaluate the impact of the regularization hyperparameter on the fulfillment of validity and probabilistic plausibility conditions, we conducted a focused hyperparameter sensitivity analysis. While theoretically should extend to infinity, practical considerations necessitate setting a feasible value. Our objective is to identify an optimal that not only guarantees condition fulfillment but also to understand its influence on other metrics. Experiments were carried out on the Moons and Law datasets, exploring values within the set . The results in Tab. 4 indicate that moderate values of , like or , deliver satisfactory outcomes, while values around or more almost invariably guarantee the fulfillment of the conditions, leading us to adopt the value of for all preceding experiments. This experiment confirms the expected trade-off: higher strictness in counterfactual conditions leads to decreased proximity metrics, requiring larger deviations from the original data point.
@@ 7 Conclusions
Acknowledgements
Patryk Wielopolski, Oleksii Furman, and Maciej Zieba’s work was supported by the National Science Centre (Poland) Grant No. 2021/43/B/ST6/02853, and Jerzy Stefanowski’s work was supported by the National Science Centre (Poland) grant No. 2023/51/B/ST6/00545. Moreover, we gratefully acknowledge Polish high-performance computing infrastructure PLGrid (HPC Center: ACK Cyfronet AGH) for providing computer facilities and support within computational grant no. PLG/2023/016636.
@@ Appendix B Datasets
In Tab. 9, we provide detailed descriptions of the datasets utilized in our study: Moons, Law222https://www.kaggle.com/datasets/danofer/law-school-admissions-bar-passage, Audit333https://archive.ics.uci.edu/dataset/475/audit+data, Heloc444https://community.fico.com/s/explainable-machine-learning-challenge, Wine555https://archive.ics.uci.edu/dataset/109/wine, Blobs and Digits666https://archive.ics.uci.edu/dataset/80/optical+recognition+of+handwritten+digits. The Moons dataset is an artificially generated set comprising two interleaving half-circles. It includes a standard deviation of Gaussian noise set at 0.01. The Law dataset originates from the Law School Admissions Council (LSAC) and is referred to in the literature as the Law School Admissions dataset [34]. For our analysis, we selected three features that exhibit the highest correlation with the target variable: entrance exam scores (LSAT), grade-point average (GPA), and first-year average grade (FYA). The Audit dataset, which encompasses comprehensive one-year non-confidential data of firms in the years 2015 to 2016, is collected from the Auditor Office of India to build a predictor for classifying suspicious firms. The Heloc dataset, initially utilized in the ’FICO xML Challenge’, consists of Home Equity Line of Credit (HELOC) applications submitted by real homeowners. This dataset comprises various numeric features that encapsulate information from the applicant’s credit report. The primary objective is to predict whether the applicant will repay their HELOC account within a two-year period. This prediction is instrumental in determining the applicant’s qualification for a line of credit. The Wine dataset comprises chemical analysis results for wines originating from the same region in Italy, produced from three distinct cultivars. This analysis quantified 13 different constituents present in each of the three wine varieties. The Blobs dataset is an artificially generated isotropic Gaussian blobs, characterized by equal variance. The Digits dataset is utilized for the optical recognition of handwritten digits. It consists of 32x32 bitmap images that are segmented into non-overlapping 4x4 blocks. Within each block, the count of ’on’ pixels is recorded, resulting in an 8x8 input matrix. Each element of this matrix is an integer between 0 and 16.
@@ Appendix D Implementation details
In the implementation of our experiments, we utilized Python [28] as the primary programming language. The core of our computational framework was PyTorch [20], a popular open-source machine learning library. A key feature of our implementation was the gradient optimization approach, designed to be executed in batches. This approach was particularly effective, allowing us to process entire test sets in a single batch. Our experiments were conducted on an M1 Apple Silicon CPU paired with 16GB of RAM. This hardware setup was more than sufficient for our experiment needs, providing enough capacity for the computational demands of our algorithm while ensuring fast processing speeds.